What is the area of the circle if XY= 10 Ft?
Answer by Jk22(389)

XY is the equation for an hyperbola

It's known that one can pass from hyperbolic to euclidean geometry via usage of complex number :

let
X=a+ib
Y=a-ib

then XY=a^2-aib+aib+b^2=a^2+b^2=10, hence it's a circle in the complex plane with radius sqrt(10)Ft.